Bonita invested R40 000 in order to have R56 000 available in 30 months' time. The yearly rate, compounded semi-annually, is

1.     7,21%.

2.     16,00%.

3.     8,00%.

4.     14,41%.

5.     13,92%.


1
Expert's answer
2021-05-18T13:54:31-0400

Investment = 40000

Amount in 30 months = 56000

Months = 30

Hal yearly terms (n) = 30/6 =5

Interest rate for 6 months = r


Therefore,

"=> 40000\times(1+r)^n = 56000"

"=> (1+r) ^5 = \frac{56000}{40000}"

"=> 1+r = 1.40^{(\frac{1}{5})}"

"=> r = 1.0696 - 1"

"=> r = 0.0696"

"=> r = 6.96\%"

Therefore, Interest rate per annum "6.96\%\times2 =13.92\%"

The correct answer is option 5